OpenStudy (anonymous):

Solve for x: x^2 -x -1= 0

OpenStudy (anonymous):

Possible intermediate steps: x^2-x-1 = 0 Add 1 to both sides: x^2-x = 1 Add 1/4 to both sides: x^2-x+1/4 = 5/4 Factor the left hand side: (x-1/2)^2 = 5/4 Take the square root of both sides: sqrt(x-1/2) = sqrt(5)/2 Eliminate the absolute value: x-1/2 = -sqrt(5)/2 or x-1/2 = sqrt(5)/2 Add 1/2 to both sides: x = 1/2 (1-sqrt(5)) or x-1/2 = sqrt(5)/2 Add 1/2 to both sides: x = 1/2 (1-sqrt(5)) or x = 1/2 (1+sqrt(5))